OpenStudy (anonymous):

If f(x) = x^2 + x and g(x) =(sqrtx-1), what is f of g(x) If f(x) = x^2 + x and g(x) =(sqrtx-1), what is f of g(x) @Mathematics

OpenStudy (anonymous):

replace \[\sqrt{-1}\] with i then you'll have \[i^2 + i\] which is a complex number with the value -1+i

OpenStudy (across):

huh?

OpenStudy (across):

\[f(x)=x^2+x\]\[g(x)=\sqrt{x-1}\]\[f(g(x))=(\sqrt{x-1})^2-(\sqrt{x-1})=x-1-\sqrt{x-1}\]

OpenStudy (across):

Change the - to a +... gosh I hate my typos...